the sum of the digit of a two digit no. is 15if the no. formed by reversing the digit is less than the original no. by 27.find the original no.

Answer by Nate(3500)

Number: 10a + b or a,b
a + b = 15 or b = 15 - a
and
10b + a = 10a + b - 27
Plug:
10b + a = 10a + b - 27
10(15 - a) + a = 10a + 15 - a - 27
150 - 10a + a = 10a + 15 - a - 27
150 - 9a = 9a - 12
162 = 18a
9 = a
b = 15 - a = 15 - 9 = 6
Number: 10a + b = 10(9) + 6 = 96